Solve the formula for the area of a triangle A = (1/2) bh for height h.
Answer by malakumar_kos@yahoo.com(315)


Solve the formula for the area of a triangle A = (1/2) bh for height h.

Area of triangle = (1/2)bh
A = (1/2)bh
2A = bh
h = 2A/b is the solution